# Mean field equation and Green’s function on the Tate curve

### 16 Mar 2026 v1

@@ 2.3 Construction of the Green’s function when $m=1$
$G(x,y)=\frac{1}{1-p^{-1}}\sum_{\phi_{h,c,n}\in\text{Char}(\mathbb{Z}_{p}^{*})\backslash\{1\}}\frac{\phi_{h,c,n}(x)\overline{\phi_{h,c,n}(y)}}{\lambda_{n}}+C$ (2.2)
Let $H_{n}$ denote the set of characters with conductor at most $n$ and define $C_{n}:=H_{n}\backslash H_{n-1}$. By the second orthogonality relation of the character, we have
$\sum_{\phi\in H_{n}}\phi(x)\overline{\phi(y)}=\boldsymbol{1}_{1+p^{n}\mathbb{Z}_{p}}(xy^{-1})|H_{n}|$ (2.3)
where $|H_{n}|=(p-1)p^{n-1}=\varphi(p^{n})$. And we notice that if $v_{p}(x-y)=M$ and $n\geq M+1$, the sum (2.3) vanishes. [^1] The idea of the finite sum formula for the case $m=1$ originated from Ekam’s independent study. Therefore,
@@
$=\frac{1}{1-p^{-1}}(\sum_{n=1}^{M}\frac{\varphi(p^{n})-\varphi(p^{n-1})}{\lambda_{n}}-\frac{\varphi(p^{M})}{\lambda_{M+1}})+C$
$=-\sum_{n=1}^{M}\frac{(p-1)p^{n}}{p^{n}+p^{n-1}-2}+\frac{p^{M+1}}{p^{M+1}+p^{M}-2}+C$
@@ 2.4 General $m$ case
$=$ $|x|_{p}\int_{\cup_{k=0}^{m-1}p^{k}\mathbb{Z}_{p}^{*}}\frac{\phi_{h,c,n}(p^{-l}z)\boldsymbol{1}_{p^{l}\mathbb{Z}_{p}^{*}}(z)-\phi_{h,c,n}(p^{-l}x)\boldsymbol{1}_{p^{l}\mathbb{Z}_{p}^{*}}(x)}{|z-x|_{p}^{2}}dz$
$=$ $|x|_{p}\int_{p^{l}\mathbb{Z}_{p}^{*}}\frac{\phi_{h,c,n}(p^{-l}z)-\phi_{h,c,n}(p^{-l}x)\boldsymbol{1}_{p^{l}\mathbb{Z}_{p}^{*}}(x)}{|z-x|_{p}^{2}}dz-\left(|x|_{p}\int_{\cup_{k\neq l}p^{k}\mathbb{Z}_{p}^{*}}\frac{1}{|z-x|_{p}^{2}}dz\right)\phi_{h,c,n}(p^{-l}x)\boldsymbol{1}_{p^{l}\mathbb{Z}_{p}^{*}}(x)$
@@
$=$ $|p^{-l}x|_{p}\int_{\mathbb{Z}_{p}^{*}}\frac{\phi_{h,c,n}(z)-\phi_{h,c,n}(p^{-l}x)\boldsymbol{1}_{p^{l}\mathbb{Z}_{p}^{*}}(x)}{|z-p^{-l}x|_{p}^{2}}dz$
when $v_{p}(x)=l$, (LABEL:S:eq_3) reduces to
$\int_{\mathbb{Z}_{p}^{*}}\frac{\phi_{h,c,n}(z)-\phi_{h,c,n}(p^{-l}x)}{|z-p^{-l}x|_{p}^{2}}dz=(D_{0}\phi_{h,c,n})(p^{-l}x)=\lambda_{n}\phi_{h,c,n}(p^{-l}x)$
@@
$-\sum_{\begin{subarray}{c}k=0\\ k\neq l\end{subarray}}^{m-1}\frac{p^{-l}}{\text{max}\{p^{-k},p^{-l}\}^{2}}p^{-k}(1-p^{-1})\phi_{h,c,n}(p^{-l}x)$
@@
$-\sum_{\begin{subarray}{c}k=0\\ k\neq l\end{subarray}}^{m-1}\frac{p^{-l}p^{-k}}{\text{max}\{p^{-k},p^{-l}\}^{2}}(1-p^{-1})$
$=$ $-(1-p^{-1})(\sum_{k=0}^{l-1}p^{-l}p^{-k}p^{2k}+\sum_{k=l+1}^{m-1}p^{-l}p^{-k}p^{2l})$
$=$ $-(1-p^{-1})(p^{-l}\sum_{k=0}^{l-1}p^{k}+\sum_{k=l+1}^{m-1}p^{-(k-l)})$
$=$ $\frac{p^{-l}+p^{l+1-m}-2}{p}$
thus (LABEL:S:eq_2) becomes
@@
$=$ $-(p^{n-1}+p^{n-2}-\frac{p^{-l}+p^{l+1-m}}{p})\phi_{h,c,n}(p^{-l}x)$
when $v_{p}(x)\neq l$, (LABEL:S:eq_3) becomes
$\frac{|p^{-l}x|_{p}}{\text{max}\{1,|p^{-l}x|_{p}\}^{2}}\int_{\mathbb{Z}_{p}^{*}}\phi_{h,c,n}(z)dz=0$
@@
$D(\phi_{h,c,n}(p^{-l}x)\boldsymbol{1}_{p^{l}\mathbb{Z}_{p}^{*}}(x))(x)=\lambda_{n,l}\phi_{h,c,n}(p^{-l}x)\boldsymbol{1}_{p^{l}\mathbb{Z}_{p}^{*}}(x)$
@@
$D\boldsymbol{1}_{p^{l}\mathbb{Z}_{p}^{*}}(x)=|p^{-l}x|_{p}\int_{\mathbb{Z}_{p}^{*}}\frac{1-\boldsymbol{1}_{p^{l}\mathbb{Z}_{p}^{*}}(x)}{|z-p^{-l}x|_{p}^{2}}dz+\frac{p^{-l}+p^{l+1-m}-2}{p}\boldsymbol{1}_{p^{l}\mathbb{Z}_{p}^{*}}(x)$
@@
$D\sum_{l=0}^{m-1}c_{l}\boldsymbol{1}_{p^{l}\mathbb{Z}_{p}^{*}}(x)=\sum_{l=0}^{m-1}c_{l}\boldsymbol{r}_{l}=(\boldsymbol{r}_{0},\boldsymbol{r}_{1},\dots,\boldsymbol{r}_{m-1})\begin{pmatrix}c_{0}\\ c_{1}\\ \vdots\\ c_{m-1}\end{pmatrix}=\boldsymbol{R}\boldsymbol{c}$
@@
$D\sum_{l=0}^{m-1}c_{l}\boldsymbol{1}_{p^{l}\mathbb{Z}_{p}^{*}}(x)=\lambda\sum_{l=0}^{m-1}c_{l}\boldsymbol{1}_{p^{l}\mathbb{Z}_{p}^{*}}(x)=\lambda\boldsymbol{c}$
@@
$\boldsymbol{R}\boldsymbol{c}=\lambda\boldsymbol{c}$
@@
$=(1-p^{-1})\delta_{h,c,n,l}^{h^{\prime},c^{\prime},n^{\prime},l^{\prime}}$
which is also orthogonal. Lastly, $\phi_{h,c,n}^{l},\phi_{0}^{l^{\prime}}>=0$ is trivial as $\phi_{h,c,n}^{l}$ is not a constant function.
For any eigenfunction $f$, we have the decomposition $f=f_{c}+f_{0}+f_{1}+\dots+f_{m-1}$, where $f_{c}$ is constant on each $p^{k}\mathbb{Z}_{p}^{*}$ and $\text{supp}f_{l}\subseteq p^{l}\mathbb{Z}_{p}^{*}$ with zero DC component (i.e. the Fourier expansion of $f_{l}$ don’t have trivial character). We can define the Fourier transform of $f_{l}$ by pulling it back to $\mathbb{Z}_{p}^{*}$ as $\widehat{f_{l}}(\xi):=\widehat{f_{l}(p^{-l}x)}(\xi)$. Then we can expand $f$ by Fourier series, and the characters are exactly the eigenfunctions. Similar to the discussion of $m=1$, this establishes completeness. ∎
@@
Then we have the Green’s function as
$G(x,y)=\frac{1}{1-p^{-1}}(\sum_{k=0}^{m-1}\sum_{\phi_{n}^{k}\in\text{Char}(p^{k}\mathbb{Z}_{p}^{*})\backslash\{1\}}\frac{\phi_{n}^{k}(x)\overline{\phi_{n}^{k}(y)}}{\lambda_{n,k}}+\sum_{k=1}^{m-1}\frac{\phi_{0}^{k}(x)\overline{\phi_{0}^{k}(y)}}{\lambda_{0,k}})+C(|y|_{p})$ (2.7)
Similarly, let $H_{n}^{l}$ denote the set of characters generated from $p^{l}\mathbb{Z}_{p}^{*}$ with conductor at most $n$ and define $C_{n}^{l}:=H_{n}^{l}\backslash H_{n-1}^{l}$. Consider $\sum_{k=0}^{m-1}\displaystyle\sum_{\phi^{k}\in H_{n}^{k}}\phi^{k}(x)\overline{\phi^{k}(y)}$. When $v_{p}(x)\neq v_{p}(y)$ (i.e. $v_{p}(d(x,y))=0$), the sum vanishes. When $v_{p}(x)=v_{p}(y)=l$ (i.e. $v_{p}(d(x,y))>0$), then
@@
$=\boldsymbol{1}_{1+p^{n}\mathbb{Z}_{p}}(xy^{-1})\varphi(p^{n})$
@@
$G(x,y)=$ $\frac{1}{1-p^{-1}}(\sum_{n=1}^{M+1}\sum_{\phi\in C_{n}^{l}}\frac{\phi(x)\overline{\phi(y)}}{\lambda_{n,l}}+\sum_{k=1}^{m-1}\frac{\phi_{0}^{k}(x)\overline{\phi_{0}^{k}(y)}}{\lambda_{0,k}})+C(|y|_{p})$ (2.8)
$=$ $\frac{1}{1-p^{-1}}(\sum_{n=1}^{M}\frac{\varphi(p^{n})-\varphi(p^{n-1})}{\lambda_{n,l}}-\frac{\varphi(p^{M})}{\lambda_{M+1,l}}+\sum_{k=1}^{m-1}\frac{q_{v_{p}(x)k}q_{v_{p}(y)k}}{\lambda_{0,k}})+C(|y|_{p})$
@@
$=\frac{1}{1-p^{-1}}((\sum_{n=1}^{M}(\varphi(p^{n})-\varphi(p^{n-1}))-\varphi(p^{M}))\delta_{v_{p}(x)}^{v_{p}(y)}+\sum_{k=1}^{m-1}q_{v_{p}(x)k}q_{v_{p}(y)k})$
$=\frac{1}{1-p^{-1}}(-\varphi(1)\delta_{v_{p}(x)}^{v_{p}(y)}+\sum_{k=0}^{m-1}q_{v_{p}(x)k}q_{v_{p}(y)k}-q_{v_{p}(x)0}q_{v_{p}(y)0})$
$=\frac{1}{1-p^{-1}}(-\delta_{v_{p}(x)}^{v_{p}(y)}+\delta_{v_{p}(x)}^{v_{p}(y)}-\frac{1}{m})$
@@
We can compute the heat kernel by using spectrum of $D$ as
$H(t,x,y)=$ $\frac{1}{1-p^{-1}}(\sum_{n=1}^{M+1}\sum_{\phi\in C_{n}^{l}}e^{\lambda_{n,l}t}\phi(x)\overline{\phi(y)}+\sum_{k=1}^{m-1}e^{\lambda_{0,k}t}\phi_{0}^{k}(x)\overline{\phi_{0}^{k}(y)})$ (2.11)
$=$ $\frac{1}{1-p^{-1}}(\sum_{n=1}^{M}e^{\lambda_{n,l}t}(\varphi(p^{n})-\varphi(p^{n-1}))-e^{\lambda_{M+1,l}t}\varphi(p^{M})+\sum_{k=1}^{m-1}e^{\lambda_{0,k}t}q_{v_{p}(x)k}q_{v_{p}(y)k})$

### 08 Apr 2026 v2

@@ 2.3 Construction of the Green’s function when $m=1$
$G(x,y)=\frac{1}{1-p^{-1}}\sum_{\phi_{h,c,n}\in\text{Char}(\mathbb{Z}_{p}^{*})\backslash\{1\}}\frac{\phi_{h,c,n}(x)\overline{\phi_{h,c,n}(y)}}{\lambda_{n}}+C.$ (2.2)
Letting $H_{n}$ denote the set of characters with conductor at most $n$ and defining $C_{n}:=H_{n}\backslash H_{n-1}$. By the second orthogonality relation of the characters, we have
$\sum_{\phi\in H_{n}}\phi(x)\overline{\phi(y)}=\boldsymbol{1}_{1+p^{n}\mathbb{Z}_{p}}(xy^{-1})|H_{n}|,$ (2.3)
where $|H_{n}|=(p-1)p^{n-1}=\varphi(p^{n})$. If $v_{p}(x-y)=M$ and $n\geq M+1$, the sum (2.3) vanishes. [^1] The idea of the finite sum formula for the case $m=1$ originated from Ekam’s independent study. Therefore,
@@
$=\frac{1}{1-p^{-1}}\left(\sum_{n=1}^{M}\frac{\varphi(p^{n})-\varphi(p^{n-1})}{\lambda_{n}}-\frac{\varphi(p^{M})}{\lambda_{M+1}}\right)+C$
$=-\sum_{n=1}^{M}\frac{(p-1)p^{n}}{p^{n}+p^{n-1}-2}+\frac{p^{M+1}}{p^{M+1}+p^{M}-2}+C,$
@@ 2.4 General $m$ case
$=$ $|x|_{p}\int_{\bigcup_{k=0}^{m-1}p^{k}\mathbb{Z}_{p}^{*}}\frac{\phi_{h,c,n}(p^{-l}z)\boldsymbol{1}_{p^{l}\mathbb{Z}_{p}^{*}}(z)-\phi_{h,c,n}(p^{-l}x)\boldsymbol{1}_{p^{l}\mathbb{Z}_{p}^{*}}(x)}{|z-x|_{p}^{2}}dz$
$=$ $|x|_{p}\int_{p^{l}\mathbb{Z}_{p}^{*}}\frac{\phi_{h,c,n}(p^{-l}z)-\phi_{h,c,n}(p^{-l}x)\boldsymbol{1}_{p^{l}\mathbb{Z}_{p}^{*}}(x)}{|z-x|_{p}^{2}}dz-\left(|x|_{p}\int_{\bigcup_{k\neq l}p^{k}\mathbb{Z}_{p}^{*}}\frac{1}{|z-x|_{p}^{2}}dz\right)\phi_{h,c,n}(p^{-l}x)\boldsymbol{1}_{p^{l}\mathbb{Z}_{p}^{*}}(x),$
@@
$=$ $|p^{-l}x|_{p}\int_{\mathbb{Z}_{p}^{*}}\frac{\phi_{h,c,n}(z)-\phi_{h,c,n}(p^{-l}x)\boldsymbol{1}_{p^{l}\mathbb{Z}_{p}^{*}}(x)}{|z-p^{-l}x|_{p}^{2}}dz,$
when $v_{p}(x)=l$, (LABEL:eq:compute_eigfunction_general_m_first_term) reduces to
$\int_{\mathbb{Z}_{p}^{*}}\frac{\phi_{h,c,n}(z)-\phi_{h,c,n}(p^{-l}x)}{|z-p^{-l}x|_{p}^{2}}dz=(D_{0}\phi_{h,c,n})(p^{-l}x)=\lambda_{n}\phi_{h,c,n}(p^{-l}x),$
@@
$-\sum_{\begin{subarray}{c}k=0\\ k\neq l\end{subarray}}^{m-1}\frac{p^{-l}}{\max\{p^{-k},p^{-l}\}^{2}}p^{-k}(1-p^{-1})\phi_{h,c,n}(p^{-l}x).$
@@
$-\sum_{\begin{subarray}{c}k=0\\ k\neq l\end{subarray}}^{m-1}\frac{p^{-l}p^{-k}}{\max\{p^{-k},p^{-l}\}^{2}}(1-p^{-1})$
$=$ $-(1-p^{-1})\left(\sum_{k=0}^{l-1}p^{-l}p^{-k}p^{2k}+\sum_{k=l+1}^{m-1}p^{-l}p^{-k}p^{2l}\right)$
$=$ $-(1-p^{-1})\left(p^{-l}\sum_{k=0}^{l-1}p^{k}+\sum_{k=l+1}^{m-1}p^{-(k-l)}\right)$
$=$ $\frac{p^{-l}+p^{l+1-m}-2}{p},$
thus (LABEL:eq:compute_eigfunction_general_m) becomes
@@
$=$ $-(p^{n-1}+p^{n-2}-\frac{p^{-l}+p^{l+1-m}}{p})\phi_{h,c,n}(p^{-l}x).$
When $v_{p}(x)\neq l$, (LABEL:eq:compute_eigfunction_general_m_first_term) becomes
$\frac{|p^{-l}x|_{p}}{\max\{1,|p^{-l}x|_{p}\}^{2}}\int_{\mathbb{Z}_{p}^{*}}\phi_{h,c,n}(z)dz=0$
@@
$D(\phi_{h,c,n}(p^{-l}x)\boldsymbol{1}_{p^{l}\mathbb{Z}_{p}^{*}}(x))(x)=\lambda_{n,l}\phi_{h,c,n}(p^{-l}x)\boldsymbol{1}_{p^{l}\mathbb{Z}_{p}^{*}}(x),$
@@
$D\boldsymbol{1}_{p^{l}\mathbb{Z}_{p}^{*}}(x)=|p^{-l}x|_{p}\int_{\mathbb{Z}_{p}^{*}}\frac{1-\boldsymbol{1}_{p^{l}\mathbb{Z}_{p}^{*}}(x)}{|z-p^{-l}x|_{p}^{2}}dz+\frac{p^{-l}+p^{l+1-m}-2}{p}\boldsymbol{1}_{p^{l}\mathbb{Z}_{p}^{*}}(x),$
@@
$D\sum_{l=0}^{m-1}c_{l}\boldsymbol{1}_{p^{l}\mathbb{Z}_{p}^{*}}(x)=\sum_{l=0}^{m-1}c_{l}\boldsymbol{r}_{l}=(\boldsymbol{r}_{0},\boldsymbol{r}_{1},\dots,\boldsymbol{r}_{m-1})\begin{pmatrix}c_{0}\\ c_{1}\\ \vdots\\ c_{m-1}\end{pmatrix}=\boldsymbol{R}\boldsymbol{c},$
@@
$D\sum_{l=0}^{m-1}c_{l}\boldsymbol{1}_{p^{l}\mathbb{Z}_{p}^{*}}(x)=\lambda\sum_{l=0}^{m-1}c_{l}\boldsymbol{1}_{p^{l}\mathbb{Z}_{p}^{*}}(x)=\lambda\boldsymbol{c},$
@@
$\boldsymbol{R}\boldsymbol{c}=\lambda\boldsymbol{c},$
@@
$\phi_{0}^{l}(x)=\boldsymbol{q}_{l}(x)=\sum_{k=0}^{m-1}q_{k,l}\boldsymbol{1}_{p^{k}\mathbb{Z}_{p}^{*}}(x)$
and
$\phi_{h,c,n}^{l}(x)=\phi_{h,c,n}(p^{-l}x)\boldsymbol{1}_{p^{l}\mathbb{Z}_{p}^{*}}(x),$
we have the following lemma:
@@
$=(1-p^{-1})\delta_{h,c,n,l}^{h^{\prime},c^{\prime},n^{\prime},l^{\prime}},$
which is also orthogonal. Lastly, $\langle\phi_{h,c,n}^{l},\phi_{0}^{l^{\prime}}\rangle=0$ is trivial as $\phi_{h,c,n}^{l}$ is not a constant function on each stratum $p^{k}\mathbb{Z}_{p}^{*}$.
For any eigenfunction $f$, we have the decomposition $f=f_{c}+f_{0}+f_{1}+\dots+f_{m-1}$, where $f_{c}$ is constant on each stratum and $\text{supp}f_{l}\subseteq p^{l}\mathbb{Z}_{p}^{*}$ with zero DC component (i.e. the Fourier expansion of $f_{l}$ don’t have trivial character). We can define the Fourier transform of $f_{l}$ by pulling it back to $\mathbb{Z}_{p}^{*}$ as $\widehat{f_{l}}(\xi):=\widehat{f_{l}(p^{-l}x)}(\xi)$. Then we can expand $f$ by Fourier series, and the characters are exactly the eigenfunctions. Similar to the discussion of $m=1$, this establishes completeness. ∎
@@
Then the Green’s function is
$G(x,y)=\frac{1}{1-p^{-1}}\left(\sum_{k=0}^{m-1}\sum_{\phi_{n}^{k}\in\text{Char}(p^{k}\mathbb{Z}_{p}^{*})\backslash\{1\}}\frac{\phi_{n}^{k}(x)\overline{\phi_{n}^{k}(y)}}{\lambda_{n,k}}+\sum_{k=1}^{m-1}\frac{\phi_{0}^{k}(x)\overline{\phi_{0}^{k}(y)}}{\lambda_{0,k}}\right)+C(|y|_{p}).$ (2.7)
Similarly, letting $H_{n}^{l}$ denote the set of characters generated from $p^{l}\mathbb{Z}_{p}^{*}$ with conductor at most $n$ and defining $C_{n}^{l}:=H_{n}^{l}\backslash H_{n-1}^{l}$. Considering
$\sum_{k=0}^{m-1}\sum_{\phi^{k}\in H_{n}^{k}}\phi^{k}(x)\overline{\phi^{k}(y)}.$
When $v_{p}(x)\neq v_{p}(y)$ (i.e. $v_{p}(d(x,y))=0$), the sum vanishes. When $v_{p}(x)=v_{p}(y)=l$ (i.e. $v_{p}(d(x,y))\geq 0$), then
@@
$=\boldsymbol{1}_{1+p^{n}\mathbb{Z}_{p}}(xy^{-1})\varphi(p^{n}).$
@@
$G(x,y)=$ $\frac{1}{1-p^{-1}}\left(\sum_{n=1}^{M+1}\sum_{\phi\in C_{n}^{l}}\frac{\phi(x)\overline{\phi(y)}}{\lambda_{n,l}}+\sum_{k=1}^{m-1}\frac{\phi_{0}^{k}(x)\overline{\phi_{0}^{k}(y)}}{\lambda_{0,k}}\right)+C(|y|_{p})$ (2.8)
$=$ $\frac{1}{1-p^{-1}}\left(\sum_{n=1}^{M}\frac{\varphi(p^{n})-\varphi(p^{n-1})}{\lambda_{n,l}}-\frac{\varphi(p^{M})}{\lambda_{M+1,l}}+\sum_{k=1}^{m-1}\frac{q_{v_{p}(x)k}q_{v_{p}(y)k}}{\lambda_{0,k}}\right)+C(|y|_{p})$
@@
$=\frac{1}{1-p^{-1}}\left(\left(\sum_{n=1}^{M}(\varphi(p^{n})-\varphi(p^{n-1}))-\varphi(p^{M})\right)\delta_{v_{p}(x)}^{v_{p}(y)}+\sum_{k=1}^{m-1}q_{v_{p}(x)k}q_{v_{p}(y)k}\right)$
$=\frac{1}{1-p^{-1}}\left(-\varphi(1)\delta_{v_{p}(x)}^{v_{p}(y)}+\sum_{k=0}^{m-1}q_{v_{p}(x)k}q_{v_{p}(y)k}-q_{v_{p}(x)0}q_{v_{p}(y)0}\right)$
$=\frac{1}{1-p^{-1}}\left(-\delta_{v_{p}(x)}^{v_{p}(y)}+\delta_{v_{p}(x)}^{v_{p}(y)}-\frac{1}{m}\right)$
@@
Furthermore, we can compute the heat kernel by using spectrum of $D$ as
$H(t,x,y)=$ $\frac{1}{1-p^{-1}}\left(\sum_{n=1}^{M+1}\sum_{\phi\in C_{n}^{l}}e^{\lambda_{n,l}t}\phi(x)\overline{\phi(y)}+\sum_{k=1}^{m-1}e^{\lambda_{0,k}t}\phi_{0}^{k}(x)\overline{\phi_{0}^{k}(y)}\right)$ (2.9)
$=$ $\frac{1}{1-p^{-1}}\left(\sum_{n=1}^{M}e^{\lambda_{n,l}t}(\varphi(p^{n})-\varphi(p^{n-1}))-e^{\lambda_{M+1,l}t}\varphi(p^{M})+\sum_{k=1}^{m-1}e^{\lambda_{0,k}t}q_{v_{p}(x)k}q_{v_{p}(y)k}\right)$
@@
Remark 2.6.
As $\mathbb{Q}_{p}^{\times}/p^{m\mathbb{Z}}\cong p^{\mathbb{Z}/m\mathbb{Z}}\times\mathbb{Z}_{p}^{*}$ is a multiplicative group, and the characters of $p^{\mathbb{Z}/m\mathbb{Z}}$ is given by
$\chi_{k}(p^{l})=e^{\frac{2\pi ik}{m}l}$
for $k,l\in\{0,1,\dots,m-1\}$. So $\forall\ x=p^{l}(a_{0}+a_{1}p+\dots)\in\mathbb{Q}_{p}^{\times}/p^{m\mathbb{Z}}$ with $a_{0}\neq 0$, we have the character
$\chi_{k,h,c,n}(x)=\chi_{k}(p^{l})\chi_{h}(a_{0})\chi_{c,n}\left(\frac{x}{p^{l}a_{0}}\right)=e^{\frac{2\pi ik}{m}l}e^{\frac{2\pi ih}{p-1}\text{ord}(a_{0})}e^{2\pi i\left\{\frac{c\ln\left(\frac{x}{p^{l}a_{0}}\right)}{p^{n}}\right\}_{p}}.$
One can check that the eigenfunctions are the linear transformation of the characters on each stratum. So the orthogonality and completeness can be easily obtained by the orthogonality of characters and Fourier transform.
Remark 2.7.
For general $s$ case, we define the operator $D^{s}$ on the Tate curve as
$D^{s}u(x)=|x|_{p}\int_{\mathbb{Q}_{p}^{\times}/q^{\mathbb{Z}}}\frac{u(z)-u(x)}{|z-x|_{p}^{1+s}}dz,$
which is also self-adjoint, negative semi-definite and preserves locally constant functions. One can check that all our computations carry over to $D^{s}$, simply by replacing $2$ with $1+s$, so that the eigenfunctions of $D^{s}$ are the same with $D$. Hence, we may use the same method to determine the spectrum of $D^{s}$, and then obtain its Green’s function and heat kernel as a finite sum formula.